OpenStudy (callisto):

Coordinate geometry question #3

OpenStudy (callisto):

I did this question by letting vertices with the knowns (x1,y1), (x2,y2), (x3,y3) and solve it. Is there any faster way to do this question?

OpenStudy (anonymous):

i can't see any other way ... but I'll keep thinking

OpenStudy (callisto):

Alight... Thanks all. Then perhaps that's the only way :(

OpenStudy (experimentx):

good old classic brute method always works for me http://www.wolframalpha.com/input/?i=%28y1%2By2%29%2F2%3D4%2C+%28y2%2By3%29%2F2%3D0%2C+%28y1%2By3%29%2F2%3D2 still there might be some elegant way of doing it.

OpenStudy (accessdenied):

i was wondering if maybe there was some transformation / resizing that you could do to those points... D:

OpenStudy (callisto):

Draw another triangle with the mid-points? :O

OpenStudy (callisto):

Sorry, question closed for a while. I have an easier question to ask....

OpenStudy (anonymous):

Three vertices \(\left(x_1,y_1\right),\left(x_2,y_2\right) \) and \(\left(x_3,y_3\right)\). \[x_1 + x_2 = 2\cdot3, \quad x_2 + x_3 = 2\cdot2,\quad x_1+x_3=2\cdot 4\]\[y_1 + y_2 = 2\cdot4,\quad y_2 + y_3 = 2\cdot0,\quad y_1+y_3=2\cdot 2\]

OpenStudy (anonymous):

\[y_2 = -y_3\]Is the key here.

OpenStudy (experimentx):

looks like @Ishaan94 and I think in the similar way. Genius think alike!!! LOL

OpenStudy (anonymous):

Lol hehe

OpenStudy (callisto):

Hmm.. so that's the only way to solve, right?

OpenStudy (experimentx):

well, good ol easy method ..! still, there might be other methods. who knows?

OpenStudy (experimentx):

I think it would be worth while to determine these vertices points in general terms => we would have our own formula

OpenStudy (callisto):

I told my friend to do it in this way as I did. But she was shocked. She couldn't believe that we need to solve the so many equations for a MC question. But she didn't know how to solve it :S

OpenStudy (anonymous):

Hmm the equations aren't that hard to solve, yeah they are boring but you can still solve it.

OpenStudy (callisto):

I know they are not ...

OpenStudy (anonymous):

Like you should solve the y co-ordinate first, it's much easier and will help you to eliminate some of the options. \[y_1 -y_3 = 8 \text{ and } y_3 = 4-y_1\] \[2y_1 = 12 \implies y_1 = 6 \implies (y_2,y_3) =(2,-2) \]

OpenStudy (anonymous):

Now you only need the options with 2 as the y co-ordinate.

OpenStudy (callisto):

Hmm.. I know how to solve them, and I told my friend it was easy to solve

OpenStudy (callisto):

The question is that is there any other to do this question...

OpenStudy (anonymous):

Solving one of the values of x_1,x_2 and x_3, will give you all the other options easily.

OpenStudy (accessdenied):

A way of looking at it that I see is to try finding the points that make a parallelogram between two midpoints and the last midpoint + a vertex.  You'd just have to do it three times for each vertex. :O